OpenStudy (anonymous):

Which expression is equivalent to (1+9x)10 A.-10-90x B.11+9x C.10+90x D.-11-9x

pooja195 (pooja195):

distribute 10*9=? 10*1=?

OpenStudy (anonymous):

90x+10

OpenStudy (anonymous):

(1+9x)(10) =(1+9x)(10) =(1)(10)+(9x)(10) =10+90x =90x+10

pooja195 (pooja195):

-_-